# Barrier relaxations of the classical and quantum optimal transport problems

### v1

@@ 2.4. Generalized Sinkhorn algorithm
$\frac{n_{1}n_{2}\varepsilon}{2}\log\left(2c^{2}+2\big{(}\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\big{)}^{2}\right)$
@@
where $(\mathbf{u}_{1},\mathbf{u}_{2})$ is a unique point in $\mathrm{D}$. The dual characterization of $\tau(C,P)$ given by (1.20) yields that $\mathbf{p}_{1}^{\top}\mathbf{u}_{1}+\mathbf{p}_{2}^{\top}\mathbf{u}_{2}\leq\tau(C,P)$. Hence, we have the following implications:
@@
$\frac{\varepsilon}{2}\log\prod_{j_{1}=j_{2}=1}^{n_{1},n_{2}}(c_{j_{1},j_{2}}-u_{j_{1},1}-u_{j_{2},2})^{2}\leq$
$\frac{\varepsilon}{2}\log\left(\frac{\sum_{j_{1}=j_{2}=1}^{n_{1},n_{2}}(c_{j_{1},j_{2}}-u_{j_{1},1}-u_{j_{2},2})^{2}}{n_{1}n_{2}}\right)^{n_{1}n_{2}}\leq$
$-\frac{\varepsilon n_{1}n_{2}\log(n_{1}n_{2})}{2}+\frac{n_{1}n_{2}\varepsilon}{2}\log 2\sum_{j_{1}=j_{2}=1}^{n_{1},n_{2}}c_{j_{1},j_{2}}^{2}+(u_{j_{1},1}+u_{j_{2},2})^{2}<$
$-\frac{\varepsilon n_{1}n_{2}\log(n_{1}n_{2})}{2}+$
$\frac{n_{1}n_{2}\varepsilon}{2}\log 2\left(\|C\|_{F}^{2}+n_{1}n_{2}\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right)^{2}\right)$
where the last inequality follows from the first inequality in (2.9). Clearly, $\|C\|_{F}^{2}\leq n_{1}n_{2}c^{2}$. Finally we deduce the inequality inequality
@@
$\frac{n_{1}n_{2}\varepsilon}{2}\log\left(2c^{2}+2\big{(}\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\big{)}^{2}\right)$
@@
    The function $f$ is strictly convex, and $f,-f^{\prime},f^{\prime\prime}$ are strictly decreasing from $\infty$ to 00 on $(-a_{n},\infty)$.
@@
    Let $a>0$ be given and consider the equation $f(x)-a=0$. It has a unique solution in $(-a_{n},\infty)$, denoted as $x(a)$, which satisfies the following inequalities:
@@
    $(x_{k},y_{k})=\begin{cases}(z_{k},y_{k-1})\textrm{ if }f(z_{k})>a,\\ (x_{k-1},z_{k})\textrm{ if }f(z_{k})<a,\\ (z_{k},z_{k})\textrm{ if }f(z_{k})=a,\textrm{ stop }.\end{cases}$
@@
We state an analogous result to Menon’s result for positive matrices [25], which is corollary of part (b) of Lemma 2.1:
@@
Set $\varepsilon=(\mathbf{1}_{n_{1}}^{\top}\mathbf{v}_{1})^{-1}$. Then $\mathbf{p}_{i}=\varepsilon\mathbf{v}_{i},i\in[2]$ are positive probability vectors. Apply part (b) of Lemma 2.1. ∎
@@
The Sinkhorn type algorithm is a coordinate descent (minimization) [24, 1, 39, 22, 14] for the minimum of the function $-\varphi(\mathbf{x},\mathbf{y})$ given in (1.17). Hence, the above Sinkhorn type algorithm converges at least linearly.
@@ 3.1. An entropic relaxation of the classical multi-partite optimal transport problem
It is a LP problem with $\prod_{k=1}^{d}n_{k}$ nonnegative variables and $1+\sum_{k=1}^{d}(n_{k}-1)$ constraints. The MPTOT was considered in [31, 32] in the context of multidimensional assignment problem, where the entires of the tensor $\mathcal{V}$ are either 00 or $1$. There is a vast literature on the MPOTP. See for example [17, 3, 39, 19, 22, 14, 15] and the references therein. The entropic relaxation of MPOTP can be stated as follows. For a given $\varepsilon>0$ consider the minimum
@@ 3.2. A barrier relaxation of the classical multi-partite optimal transport problem
Let $(\mathbf{p}_{1},\cdots,\mathbf{p}_{d})\in(\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{d}})_{++,b}$. Assume that $\mathrm{V}(P)$ be defined by (3.1).
@@
    Consdier the minimum problem (3.8). Then
@@
    The interior of the ball $\mathrm{B}((\mathbf{x}_{1,0},\ldots,\mathbf{x}_{d,0}),r)=\{\sum_{i=1}^{d}\|\mathbf{x}_{i}-\mathbf{x}_{i.0}\|^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((\mathbf{x}_{0},\ldots,\mathbf{x}_{d,0}),r)$, contains $(\mathbf{x}_{1},\ldots,\mathbf{x}_{d})$ that satisfies the conditoins (3.15).
@@
    (3.19) $\hat{\beta}(\mathbf{x},\mathbf{y})=-\sum_{j_{1}=\cdots=j_{d}=1}^{n}\log(c_{j_{1},\cdots,j_{d}}-x_{j_{1},1}-\cdots-x_{j_{d},d})$
    $-\log\left(r^{2}-\sum_{i=1}^{d}\|\mathbf{x}-\mathbf{x}_{0,i}\|^{2}\right)+\log r^{2}$
@@
Let $d\geq 3$, and $\mathbf{p}_{i}\in\Delta^{n},i\in[d]$ be positive probability vectors. Consider the minimum (3.8), which corresponds to the maximum dual problem (3.14). Apply the IPM short step interior path algorithm with the barrier (3.19) to the $-\min$ problem as in (2.2), where $r$ is given by (3.18). Let $c=\max_{j_{i}\in[n_{i}],i\in[d]}|c_{j_{1},\ldots,j_{d}}|$. Choose a starting point given by (3.17). The CPA algorithm finds the value $\tau(C,P)$ within precision $\delta>0$ in $O\big{(}n^{d/2}\log\frac{n^{d}\sqrt{d}r}{\delta}\big{)}$ iterations. The arithmetic complexity of CPA is $O\big{(}dn^{3d/2}\log\frac{n^{d}\sqrt{d}r}{\delta}\big{)}$.
@@ 4.1. A barrier relaxation of the quantum bi-partite optimal transport problem
Let $\rho_{i}\in\mathrm{H}_{n_{i},++},n_{i}>1,i\in[2]$ and assume that $\Gamma(R)$ be defined by (1.6) for $d=2$.
@@
(a) As $-\log\det\rho$ is strictly convex in $\Gamma_{o}(R)$ and is $\infty$ on $\partial\Gamma(R)$, it follows that $\kappa_{\beta}(C,R)$ as achieved at the unique $\rho^{\star}$. Clearly $\operatorname{Tr}C\rho^{\star}\geq\kappa(C,R)$. Since $\rho\in\Gamma_{o}(R)$ we have that $\sum_{k=1}^{n_{1}n_{2}}\lambda_{k}(\rho^{\star})=1$ we deduce $10<\det\rho^{\star}<$ 1. Hence, $\beta(\rho^{\star})>0$. As $\varepsilon>0$ we deduce the inequality (4.2).
@@
$\varepsilon((\rho^{\star})^{-1})_{(i_{1},j_{1})(i_{2},j_{2})}=\begin{cases}c_{(i_{1},j_{1})(i_{2},j_{2})}\textrm{ if }i_{1}\neq i_{2}\textrm{ and }j_{1}\neq j_{2},\\ c_{(i,j_{1})(i,j_{2})}-\mu_{j_{1},j_{2}}\textrm{ if }j_{1}\neq j_{2},\\ c_{(i_{1},j)(i_{2},j)}-\lambda_{i_{1},i_{2}}\textrm{ if }i_{1}\neq i_{2},\\ c_{(i,j)(i,j)}-\lambda_{i,i}-\mu_{j,j}\end{cases}$
@@
Proof of Theorem 1.2 . As $\log\det Z$ is striclty convex of $\mathrm{H}_{(n_{1},n_{2}),++}$ it follows that $\varphi(X,Y,C)$ is concave on $\hat{\mathrm{D}}$. To show that $\varphi$ is strictly concave on $\mathrm{D}$ we need to show that if $\mathrm{C}(X,Y)=\mathrm{C}(X_{1},Y_{1})$ for $(X,Y),(X_{1},Y_{1})\in\mathrm{D}$, then $(X,Y)=(X_{1},Y_{1})$. This is equivalent to the implication (4.4), which is shown in the proof of Lemma 4.1.
@@
It is left to show that $\varphi(U_{1},U_{2},C)=\kappa_{\beta}(R,C,\varepsilon)$, which is equivalent to (1.28).
Finally, observe
@@
$\operatorname{Tr}\big{(}C(U_{1},U_{2})+U_{1}\otimes I_{n_{2}}+I_{n_{1}}\otimes U_{2})\rho^{\star}-\varepsilon\log\det\big{(}\varepsilon C(U_{1},U_{2})^{-1}\big{)}=$
$\operatorname{Tr}\big{(}C(U_{1},U_{2})\rho^{\star}+\operatorname{Tr}(U_{1}\otimes I_{n_{2}}+I_{n_{1}}\otimes U_{2})\rho^{\star}-\varepsilon\log(\varepsilon^{n_{1}n_{2}}(\det C(U_{1},U_{2})^{-1}))$
$=\varepsilon\operatorname{Tr}I_{n_{1}}\otimes I_{n_{2}}+\operatorname{Tr}U_{1}\rho_{1}+\operatorname{Tr}U_{2}\rho_{2}-n_{1}n_{2}\varepsilon\log\varepsilon+\varepsilon\log\det C(U_{1},U_{2})=$
@@ 4.2. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    (4.11) $\|(X,Y)\|_{F}=\sqrt{\|X\|_{F}^{2}+\|Y\|_{F}^{2}}<\sqrt{\max(n_{1},n_{2})}\frac{2\|C\|_{2}}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})}.$
@@
    The interior of the ball $\mathrm{B}((X_{0},Y_{0}),r)=\{\|X-X_{0}\|_{F}^{2}+\|Y-Y_{0}\|_{F}^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((X_{0},Y_{0}),r)$, contains $(X,Y)$ that satisfies the conditoins (4.10). In particular, the three supremum of (4.9) are achieved.
@@
(b) The first equality in (4.9) is proved in [9, Theorem 3.2]. The second equality is straightforward. Clearly $C(X,Y)=C(X+tI_{n_{1}},Y-tI_{n_{2}})$. As in the proof of Lemma 4.1 choose $t=\frac{\operatorname{Tr}U_{2}-\operatorname{Tr}U_{1}}{n_{1}+n_{2}}$ to deduce the third equality.
@@
Assume that $X=\operatorname{diag}(\mathbf{x}),Y=\operatorname{diag}(\mathbf{y})$. Denote by $\operatorname{diag}(C)\in\mathrm{H}_{(n_{1},n_{2})}$ the diagonal matrix obtained from $C$ by deleting the off-diagonal entries of $C$. Name the diagonal entry of $\operatorname{diag}(C)$ in the row and column $(j_{1},j_{2})$ by $c_{j_{1},j_{2}}$. Recall the well known inequalities following from the Cauchy interlacing inequalities [13, Section 4.4, Problem (3a)]:
@@
$\Rightarrow c:=\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|\leq\|C\|_{2}.$
Assume that the conditions of the implication in (4.10) hold for $X=\operatorname{diag}(\mathbf{x}),Y=\operatorname{diag}(\mathbf{y})$ hold. Thus, we have the conditions in the implication (4.10), where $C=[c_{j_{1},j_{2}}]$ and the coordinates of $\mathbf{p}_{1},\mathbf{p}_{2}$ are the diagonal elements of $\rho_{1},\rho_{2}$ respectively. Therefore, we deduce the inequalities (2.9). The Cauchy interlacing property yields that $\min_{j_{i}\in[n_{i}]}\mathbf{p}_{i}\geq\lambda_{\min}(\rho_{i})$ for $i\in[2]$. Combine the above inequalities to deduce (4.12) in this case.
Assume the conditions (4.10). Assume that $X=Q_{X}\operatorname{diag}(\lambda(X))Q_{X}^{*},Y=Q_{Y}\operatorname{diag}(\lambda(X))Q_{Y}^{*}$, where $Q_{X},Q_{Y}$ are corresponding unitary matrices. Denote
@@
(f) Observe that $(X,Y)$ is on the boundary of $D_{1}$ if either $\|X-X_{0}\|_{F}^{2}+\|Y-Y_{0}\|_{F}^{2}=r^{2}$ or $\lambda_{\min}(C(X,Y))=0$. As $r>1$ we it is enough to consider the case $\lambda_{\min}(C(X,Y))=0$. Use the Weyl’s inequalities [13, Theorem 4.4.6] and (4.16) to deduce
@@
$\frac{n_{1}n_{2}\varepsilon}{2}\log\left(2\|C\|_{2}^{2}+2\big{(}\frac{2\|C\|_{2}}{\lambda_{n_{1}}(\rho_{1})\lambda_{n_{2}}(\rho_{2})}\big{)}^{2}\right)$
@@
By conjugating $U_{1},U_{2},C$ by approriate unitary matrices, we can assume, as in the proof part (c) of Lemma 4.2, $U_{1}=\operatorname{diag}(\mathbf{x}),U_{2}=\operatorname{diag}(\mathbf{y})$ The maximal characterization of $\kappa(C,R)$ given by (4.9) yields that $\operatorname{Tr}\rho_{1}U_{1}+\operatorname{Tr}\rho_{2}U_{2}\leq\kappa(C,R)$. As $C(U_{1},U_{2})$ is a positive definite matrix it follows that $\det C(U_{1},U_{2})$ is at most the product of the diagonal entires of $C(U_{1},U_{2})$. (This is Hadamard’s determinant inequality.) As the proof of part (c) of Lemma 4.2 denote the diagonal entries of $C$ in the row and column $(j_{1},j_{2})$ by $c_{j_{1},j_{2}}$. Let $c=\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|$. Now use the arithmetic-geometric mean to deduce
$\log\det C(U_{1},U_{2})\leq\frac{1}{2}\log\prod_{j_{1}=j_{2}}^{n_{1},n_{2}}\big{(}c_{j_{1},j_{2}}-(x_{j_{1}}+y_{j_{2}})\big{)}^{2}\leq$
$\frac{1}{2}\log\left(\frac{\sum_{j_{1}=j_{2}}^{n_{1},n_{2}}\big{(}c_{j_{1},j_{2}}-(x_{j_{1}}+y_{j_{2}})\big{)}^{2}}{n_{1}n_{2}}\right)^{n_{1}n_{2}}\leq$
$\frac{1}{2}\log\left(\frac{2\sum_{j_{1}=j_{2}}^{n_{1},n_{2}}c_{j_{1},j_{2}}^{2}+(x_{j_{1}}+y_{j_{2}})^{2}}{n_{1}n_{2}}\right)^{n_{1}n_{2}}\leq$
$\frac{1}{2}\log\left(\frac{2n_{1}n_{2}\left(c^{2}+\big{(}\frac{2c}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})}\big{)}^{2}\right)}{n_{1}n_{2}}\right)^{n_{1}n_{2}}.$
For the last inequalites we used the arguments in the proof of part (c) of Lemma 4.2. Recall that $c\leq\|C\|_{2}$. This establishes (4.17). ∎

### v2

@@ 2.4. Generalized Sinkhorn algorithm
$n_{1}n_{2}\varepsilon\log\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right).$
@@
where $(\mathbf{u}_{1},\mathbf{u}_{2})$ is a unique point in $\mathrm{D}$. The dual characterization of $\tau(C,P)$ given by (1.20) yields that $\mathbf{p}_{1}^{\top}\mathbf{u}_{1}+\mathbf{p}_{2}^{\top}\mathbf{u}_{2}\leq\tau(C,P)$. Use the first set of inequalities in (2.9) to deduce
@@
$\varepsilon\log\prod_{j_{1}=j_{2}=1}^{n_{1},n_{2}}(c_{j_{1},j_{2}}-u_{j_{1},1}-u_{j_{2},2})\leq$
$\varepsilon n_{1}n_{2}\log\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right).$
Finally, we deduce the inequality
@@
$n_{1}n_{2}\varepsilon\log\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right)$
@@
    The function $f$ is strictly convex, and $f,-f^{\prime},f^{\prime\prime}$ are strictly decreasing from $\infty$ to 0 on $(-a_{n},\infty)$.
@@
    Let $a>0$ be given and consider the equation f(x)a=0f(x)-a=0. It has a unique solution in $(-a_{n},\infty)$, denoted as $x(a)$, which satisfies the following inequalities:
@@
[…]
@@
We state an analogous result to Menon’s result for positive matrices [25], which is corollary of part (b) of Lemma LABEL:minucbpotplem:
@@
Set $\varepsilon=(\mathbf{1}_{n_{1}}^{\top}\mathbf{d}_{1})^{-1}$. Then $\mathbf{p}_{i}=\varepsilon\mathbf{d}_{i},i\in[2]$ are positive probability vectors. Apply part (b) of Lemma 2.2. ∎
@@
The Sinkhorn type algorithm is a coordinate descent (minimization) [24, 1, 39, 22, 14] for the minimum of the function $\varphi(\mathbf{x},\mathbf{y})$ given in (1.17). Hence, the above Sinkhorn type algorithm converges at least linearly.
@@ 3.1. An entropic relaxation of the classical multi-partite optimal transport problem
It is a LP problem with $\prod_{k=1}^{d}n_{k}$ nonnegative variables and $1+\sum_{k=1}^{d}(n_{k}-1)$ constraints. The MPTOT was considered in [31, 32] in the context of multidimensional assignment problem, where the entires of the tensor $\mathcal{V}$ are either 0 or $1$. There is a vast literature on the MPOTP. See for example [17, 3, 39, 19, 22, 14, 15] and the references therein. The entropic relaxation of MPOTP can be stated as follows. For a given $\varepsilon>0$ consider the minimum
@@ 3.2. A barrier relaxation of the classical multi-partite optimal transport problem
Let $\mathbf{p}_{i}\in\Delta^{n_{i}},i\in[d]$. Assume that $\mathrm{V}(P)$ be defined by (3.1).
@@
    Consider the minimum problem (3.8). Then
@@
    The interior of the ball $\mathrm{B}((\mathbf{x}_{1,0},\ldots,\mathbf{x}_{d,0}),r)=\{\sum_{i=1}^{d}\|\mathbf{x}_{i}-\mathbf{x}_{i.0}\|^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((\mathbf{x}_{1,0},\ldots,\mathbf{x}_{d,0}),r)$, contains $(\mathbf{x}_{1},\ldots,\mathbf{x}_{d})$ that satisfies the conditoins (3.15).
@@
    (3.19) $\hat{\beta}(\mathbf{x}_{1},\ldots,\mathbf{x}_{d})=-\sum_{j_{1}=\cdots=j_{d}=1}^{n}\log(c_{j_{1},\cdots,j_{d}}-x_{j_{1},1}-\cdots-x_{j_{d},d})$
    $-\log\left(r^{2}-\sum_{i=1}^{d}\|\mathbf{x}_{i}-\mathbf{x}_{0,i}\|^{2}\right)+\log r^{2}$
@@
Let $d\geq 3$, and $\mathbf{p}_{i}\in\Delta^{n},i\in[d]$ be positive probability vectors. Consider the minimum (3.8), which corresponds to the maximum dual problem (3.14). Apply the IPM short step interior path algorithm with the barrier (3.19) to the $\min$ problem as in (2.2), where $r$ is given by (3.18). Let $c=\max_{j_{i}\in[n_{i}],i\in[d]}|c_{j_{1},\ldots,j_{d}}|$. Choose a starting point given by (3.17). The CPA algorithm finds the value $\tau(C,P)$ within precision $\delta>0$ in $O\big(n^{d/2}\log\frac{n^{d}\sqrt{d}r}{\delta}\big)$ iterations. The arithmetic complexity of CPA is $O\big(dn^{3d/2}\log\frac{n^{d}\sqrt{d}r}{\delta}\big)$.
@@ 4.1. A barrier relaxation of the quantum bi-partite optimal transport problem
Let $\rho_{i}\in\mathrm{H}_{n_{i},++,1},n_{i}>1,i\in[2]$ and assume that $\Gamma(R)$ be defined by (1.6) for $d=2$.
@@
(a) As $\log\det\rho$ is strictly convex in $\Gamma_{o}(R)$ and is $\infty$ on $\partial\Gamma(R)$, it follows that $\kappa_{\beta}(C,R)$ as achieved at the unique $\rho^{\star}$. Clearly $\operatorname{Tr}C\rho^{\star}\geq\kappa(C,R)$. Since $\rho^{\star}\in\Gamma_{o}(R)$ we have that $\sum_{k=1}^{n_{1}n_{2}}\lambda_{k}(\rho^{\star})=1$ we deduce $0<\det\rho^{\star}<1$. Hence, $\beta(\rho^{\star})>0$. As $\varepsilon>0$ we obtain the inequality (4.2).
@@
[…]
@@
Proof of Theorem 1.2 . As $\log\det Z$ is striclty concave on $\mathrm{H}_{(n_{1},n_{2}),++}$ it follows that $\varphi(X,Y,C)$ is concave on $\hat{\mathrm{D}}$. To show that $\varphi$ is strictly concave on $\mathrm{D}$ we need to show that if $\mathrm{C}(X,Y)=\mathrm{C}(X_{1},Y_{1})$ for $(X,Y),(X_{1},Y_{1})\in\mathrm{D}$, then $(X,Y)=(X_{1},Y_{1})$. This is equivalent to the implication (4.4), which is shown in the proof of Lemma 4.1.
@@
It is left to show that $\varphi(U_{1},U_{2},C)=\kappa_{\beta}(R,C,\varepsilon)$, which is equivalent to (1.28). Finally, observe
@@
$\operatorname{Tr}\big(C(U_{1},U_{2})+U_{1}\otimes I_{n_{2}}+I_{n_{1}}\otimes U_{2})\rho^{\star}$-
$\varepsilon\log\det\big(\varepsilon C(U_{1},U_{2})^{-1}\big)=$
$\operatorname{Tr}\big(C(U_{1},U_{2})\rho^{\star}+\operatorname{Tr}(U_{1}\otimes I_{n_{2}}+I_{n_{1}}\otimes U_{2})\rho^{\star}$-
$\varepsilon\log(\varepsilon^{n_{1}n_{2}}(\det C(U_{1},U_{2})^{-1}))$
$=\varepsilon\operatorname{Tr}I_{n_{1}}\otimes I_{n_{2}}+\operatorname{Tr}U_{1}\rho_{1}+\operatorname{Tr}U_{2}\rho_{2}$-
$n_{1}n_{2}\varepsilon\log\varepsilon+\varepsilon\log\det C(U_{1},U_{2})=$
@@ 4.2. The interior point method for finding a barrier approximation of $\kappa(C,R)$
    (4.11) $\|(X,Y)\|_{F}=\sqrt{\|X\|_{F}^{2}+\|Y\|_{F}^{2}}<\sqrt{\max(n_{1},n_{2})}\frac{2\|C\|_{2}}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})},$
    $-\frac{2\|C\|_{2}}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})}\leq x_{k}+y_{l}\leq\|\mathbb{C}\|_{2}.$
@@
    The interior of the ball $\mathrm{B}((X_{0},Y_{0}),r)=\{\|X-X_{0}\|_{F}^{2}+\|Y-Y_{0}\|_{F}^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((X_{0},Y_{0}),r)$, contains all $(X,Y)$ that satisfies the conditoins (4.10). In particular, the three suprema of (4.9) are achieved.
@@
(b) The first equality in (4.9) is proved in [9, Theorem 3.2]. The second equality is straightforward. Clearly, $C(X,Y)=C(X+tI_{n_{1}},Y-tI_{n_{2}})$. As in the proof of Lemma 4.1 choose $t=\frac{\operatorname{Tr}U_{2}-\operatorname{Tr}U_{1}}{n_{1}+n_{2}}$ to deduce the third equality.
@@
Assume that $X=\operatorname{diag}(\mathbf{x}),Y=\operatorname{diag}(\mathbf{y})$. Denote by $\operatorname{diag}(C)\in\mathrm{H}_{(n_{1},n_{2})}$ the diagonal matrix obtained from $C$ by deleting the off-diagonal entries of $C$. Name the diagonal entry of $\operatorname{diag}(C)$ in the row and column $(j_{1},j_{2})$ by $c_{j_{1},j_{2}}$, and let $\hat{C}=[c_{j_{1},j_{2}}]\in\mathbb{R}^{n_{1}\times n_{2}}$. Recall the well known inequalities following from the Cauchy interlacing inequalities [13, Section 4.4, Problem (3a)]:
@@
$\Rightarrow c=\|C\|_{2}\geq\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|.$
Assume that the conditions of the implication in (4.10) hold for $X=\operatorname{diag}(\mathbf{x}),Y=\operatorname{diag}(\mathbf{y})$ hold. Restricting these conditions to the diagonal entries of $C(X,Y)$ we obtain the conditions:
$c_{j_{1},j_{2}}-x_{j_{1}}-y_{j_{2}}\geq 0,j_{1}\in[n_{1}],j_{2}\in[n_{2}],\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}\geq\lambda_{min}(C),\mathbf{1}_{n_{1}}^{\top}\mathbf{x}=\mathbf{1}_{n_{2}}^{\top}\mathbf{y},$
where $\mathbf{p}_{i}$ are the diagonal entries of $\rho_{i}$ for $i\in[2]$. Observe that the first set of the above inequalities yield $\operatorname{Tr}\hat{C}^{\top}\mathbf{p}_{1}\mathbf{p}_{2}^{\top}\geq\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}\geq\lambda_{min}(C)$. Hence, the following inequaltiy holds:
$\lambda_{\min}(C)\leq\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}=\sum_{j_{1}=j_{2}=1}^{n_{1},n_{2}}p_{j_{1},1}p_{j_{2},2}(x_{j_{1}}+y_{j_{2}})\leq$
$p_{1,k}p_{2,l}(x_{k}+y_{l})+\sum_{(j_{1},j_{2})\in[n_{1}]\times[n_{2}]\setminus\{(k,l)\}}p_{j_{1},1}p_{j_{2},2}c_{j_{1},j_{2}}=$
$p_{1,k}p_{2,l}(x_{k}+y_{l})-p_{1,k}p_{2,l}c_{k,l}+\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top}).$
Combine above conditions and the inequality to obtain:
(4.16) $0\leq c_{k,l}-x_{k}-y_{l}\leq\frac{\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top})-\lambda_{\min}(C)}{p_{1,k}p_{2,l}}\leq$
$\frac{\lambda_{\max}(C)-\lambda_{min}(C)}{p_{1,k}p_{2,l}}\leq\frac{2c}{p_{1,k}p_{2,l}}\leq\frac{2c}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})},$
$-\frac{2c}{p_{1,k}p_{2,l}}<\frac{c_{k,l}p_{1,k}p_{2,l}-\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top})+\lambda_{min}(C)}{p_{1,k}p_{2,l}}\leq x_{k}+y_{l}.$
The Cauchy interlacing property yields that $\min_{j_{i}\in[n_{i}]}\mathbf{p}_{i}\geq\lambda_{\min}(\rho_{i})$ for $i\in[2]$. This proves the above inequalities. Combine the above inequalities to deduce (4.11) in this case.
Assume the conditions (4.10). Let
$X=Q_{X}\operatorname{diag}(\lambda(X))Q_{X}^{*},\quad Y=Q_{Y}\operatorname{diag}(\lambda(X))Q_{Y}^{*},$
where $Q_{X},Q_{Y}$ are corresponding unitary matrices. Denote
@@
(f) Observe that $(X,Y)$ is on the boundary of $D_{1}$ if either $\|X-X_{0}\|_{F}^{2}+\|Y-Y_{0}\|_{F}^{2}=r^{2}$ or $\lambda_{\min}(C(X,Y))=0$. As $r>1$ it is enough to consider the case $\lambda_{\min}(C(X,Y))=0$. Use the Weyl’s inequalities [13, Theorem 4.4.6] and (4.17) to deduce
@@
$n_{1}n_{2}\varepsilon\log\left(\frac{2\|C\|_{2}}{\lambda_{n_{1}}(\rho_{1})\lambda_{n_{2}}(\rho_{2})}\right)$
@@
By conjugating $U_{1},U_{2},C$ by approriate unitary matrices, we can assume, as in the proof part (c) of Lemma 4.2, $U_{1}=\operatorname{diag}(\mathbf{x}),U_{2}=\operatorname{diag}(\mathbf{y})$ The maximal characterization of $\kappa(C,R)$ given by (4.9) yields that $\operatorname{Tr}\rho_{1}U_{1}+\operatorname{Tr}\rho_{2}U_{2}\leq\kappa(C,R)$. As $C(U_{1},U_{2})$ is a positive definite matrix it follows that $\det C(U_{1},U_{2})$ is at most the product of the diagonal entires of $C(U_{1},U_{2})$. (This is Hadamard’s determinant inequality.) As the proof of part (c) of Lemma 4.2 denote the diagonal entries of $C$ in the row and column $(j_{1},j_{2})$ by $c_{j_{1},j_{2}}$. Let $c=\|C\|_{2}$. Use the inequalities (4.16) to obtain
$\log\det C(U_{1},U_{2})\leq\log\prod_{j_{1}=j_{2}}^{n_{1},n_{2}}\big(c_{j_{1},j_{2}}-(x_{j_{1}}+y_{j_{2}})\leq$
$\log\big(\frac{2\|C\|_{2}}{\lambda_{\min}(\rho_{1})\lambda_{\min}(\rho_{2})}\big)^{n_{1}n_{2}}.$
This establishes (4.18). ∎